OpenStudy (anonymous):

If x^2 + 2xy + y^2 = 81, x + y= (a) 7 (b) 8 (c) 9 (d) 10 (e) 11

OpenStudy (alexwee123):

x^2 + 2xy + y^2 = (x+y)^2 so... (x+y)^2=81 x+y=9

OpenStudy (alexwee123):

so C would be your answer ^.^

OpenStudy (anonymous):

\[(x+y)^2 = x^2 + y^2 + 2xy \implies (x+y)^2 = 81 \implies x+y = 9\]

OpenStudy (anonymous):

thank you guys

OpenStudy (alexwee123):

any time ^.^

OpenStudy (anonymous):

Welcome..